OpenStudy (anonymous):

Why does the integral from -infinity to infinity of x/(x^2 +1)^2 diverge?

OpenStudy (anonymous):

Are you sure it diverges? the function\[\frac{x}{(x^2+1)^2}\]is an odd function, so any integral:\[\int\limits_{-n}^{n}\frac{x}{(x^2+1)^2}dx=0\] as n goes to infinity, that shouldnt change.

OpenStudy (anonymous):

you forgot to square the denominator >.<

OpenStudy (anonymous):

You're right, my mistake! -.-

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+x%2F%28%28x%5E2+%2B+1%29%5E2%29+ Fixed and convergent :-)

OpenStudy (anonymous):

believe in your answers :)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+from+-infinity+to+infinity+%28x%5E2+%2B1%29%5E-2+*+x Wrong link, haha :-) And, yes, @joemath314159 is correct. If you are sure about your answer, believe it :-)